# Ramification filtration via deformations, II

### v3

@@ 1. Lie algebra $\mathcal{L}$ and its filtration
As earlier, $\mathcal{K}$ is a complete discrete valuation field of characteristic $p$ with a finite residue field $k\simeq\mathbb{F}_{p^{N_{0}}}$ and a fixed uniformiser $t_{0}$, i.e. $\mathcal{K}=k((t_{0}))$. Denote by $\mathcal{K}_{sep}$ a separable closure of $\mathcal{K}$. Let $\mathcal{G}=\mathrm{Gal}(\mathcal{K}_{sep}/\mathcal{K})$. Then $\mathcal{G}$ acts on $\mathcal{K}_{sep}$ and for any $a\in\mathcal{K}_{sep}$ and $g_{1},g_{2}\in\mathcal{G}$, $g_{1}(g_{2}a)=(g_{1}g_{2})a$. Denote by $\sigma$ the $p$-th power map on $\mathcal{K}$, $\mathcal{K}_{sep}$ and $k$.
@@ 1.1. Lie algebra $\mathcal{L}$
For any field extension $\mathcal{E}$ of $\mathcal{K}$ in $\mathcal{K}_{sep}$, we have a lift modulo $p^{M}$, $O_{M}(\mathcal{E}):=W_{M}(\sigma^{M-1}\mathcal{E})[t]\subset W_{M}(\mathcal{E})$, where $t=[t_{0}]\in W_{M}(\mathcal{K})$ is the Teichmuller representative of $t_{0}$, cf. e.g. [2]. The system of lifts $O_{M}(\mathcal{E})$ is functorial in $\mathcal{E}$, the actions of $\mathcal{G}$ and $\sigma$ can be naturally extended to $O_{M}(\mathcal{K}_{sep})$. Note that $O_{M}(\mathcal{K}_{sep})^{\mathcal{G}}=O_{M}(\mathcal{K})=W_{M}(k)((t))$ and $O_{M}(\mathcal{K}_{sep})|_{\sigma=\mathrm{id}}=W_{M}(\mathbb{F}_{p})$.
Let $\mathbb{Z}^{+}(p)=\{a\in\mathbb{N}\ |\ \mathrm{gcd}\,(a,p)=1\}$, $\mathbb{Z}^{0}(p)=\mathbb{Z}^{+}(p)\cup\{0\}$.
Let ^$\widehat{\mathcal{L}}_{W_{M}(k)}$ be a profinite free Lie $W_{M}(k)$-algebra with the (topological) module of (free) generators
@@
Define $\sigma$-action on ^$\widehat{\mathcal{L}}_{W_{M}(k)}$ by setting $\sigma:D_{an}\mapsto D_{a,n+1}$ and $\sigma:D_{0}\mapsto D_{0}$. Then ^$\widehat{\mathcal{L}}_{W_{M}(k)}|_{\sigma=\mathrm{id}}:=\widehat{\mathcal{L}}$ is a free profinite Lie $\mathbb{Z}_{p}/p^{M}$-algebra such that ^$\widehat{\mathcal{L}}\otimes W_{M}(k)=\widehat{\mathcal{L}}_{W_{M}(k)}$. We set :=^$\mathcal{L}:=\widehat{\mathcal{L}}/C_{p}(\widehat{\mathcal{L}})$. This is a maximal quotient of $\widehat{\mathcal{L}}$ of nilpotent class $<p$.
@@ 1.2. Ideals $\mathcal{L}(w)$, $w\geqslant 1$
Definition.
For $w\geqslant 1$, let $\mathcal{L}(w)$ be the minimal central filtration of ideals in $\mathcal{L}$ such that if $a/v_{0}\geqslant U(w-1,m)$ then $p^{m}D_{an}\in\mathcal{L}(w)_{W_{M}(k)}$.
Note that $\mathcal{L}(1)=\mathcal{L}$ and $\forall w_{1},w_{2}\in\mathbb{N}$, $[\mathcal{L}(w_{1}),\mathcal{L}(w_{2})]\subset\mathcal{L}(w_{1}+w_{2})$.
@@
b) For $s\geqslant 1$, let $m=su_{0}+s_{1}$, where $0\leqslant s_{1}<s$. Then
$U(s,m)=s_{0}/p^{u_{0}}+s_{1}/p^{u_{0}+1}\,\text{with }s_{0}=s-s_{1}\,.$
In particular, $u_{0},s_{0},s_{1}$ are uniquely determined by $m$ and $s$.
c) Suppose $s+s^{\prime}<p$. Then for any $M\geqslant 0$,
$\min\{U(s,m)+U(s^{\prime},m^{\prime})\ |\ m+m^{\prime}=M\}=U(s+s^{\prime},M)\,.$
@@
Therefore, the minimal value $U(s,m)$ can appear only if there is $u_{0}\in\mathbb{N}$ such that among all numbers $m_{i}$, $1\leqslant j\leqslant s$, there are $s_{0}>0$ of those which are equal to $u_{0}$ and all remaining $s_{1}=s-s_{0}$ numbers are equal to $u_{0}+1$. It remains to note that $m=m_{1}+\dots+m_{s}=u_{0}s_{0}+(u_{0}+1)s_{1}=su_{0}+s_{1}\,.$
c) Suppose $U(s,m)+U(s^{\prime},m^{\prime})$ is the minimal value of the LHS. We can assume that $s,s^{\prime}\geqslant 1$. Let $m=si+s_{1}$ and $m^{\prime}=s^{\prime}i^{\prime}+s_{1}^{\prime}$, where $0\leqslant s_{1}<s$ and $0\leqslant s_{1}^{\prime}<s^{\prime}$. Then for $s_{0}=s-s_{1}$ and $s_{0}^{\prime}=s^{\prime}-s_{1}^{\prime}$, $s_{0}+s_{1}+s_{0}^{\prime}+s_{1}^{\prime}<p$, $U(s,m)=s_{0}/p^{i}+s_{1}/p^{i+1}$ and $U(s^{\prime},m^{\prime})=s_{0}^{\prime}/p^{i^{\prime}}+s_{1}^{\prime}/p^{i^{\prime}+1}$.
@@
Example. Suppose $m=0$ and $s\geqslant 1$. Then $u_{0}=s_{1}=0$ and for all $w\geqslant 1$, $U(w-1,0)=w-1$. If $(w-1)v_{0}\leqslant a<wv_{0}$ then $D_{an}\in\mathcal{L}(w)_{W_{M}(k)}$. In particular, if $a\geqslant(p-1)v_{0}$ then $D_{an}\in\mathcal{L}(p)_{W_{M}(k)}$. Notice that in the modulo $p$ situation (i.e. when $M=1$) from [12] the ideals $\mathcal{L}(w)$ coincide with the ideals introduced in that paper (and denoted by $\mathcal{L}(s)$).
Proposition 1.1b) means:
@@
The filtration $\{\mathcal{L}(w)\ |\ w\geqslant 1\}$ is the minimal central filtration of $\mathcal{L}$ such that for all $a\in\mathbb{Z}^{0}(p)$ and non-negative integers $\alpha_{0}\neq 0,\alpha_{1},j$, hold the conditions
@@
$p^{\alpha_{0}j+\alpha_{1}(j+1)}D_{an}\in\mathcal{L}(1+\alpha_{0}+\alpha_{1})_{W_{M}(k)}\,.$
1.3. Independent relations $p^{M(a,s)}D_{an}\in\mathcal{L}(1+s)_{W_{M}(k)}$, $s<p$
@@
Given $a\in\mathbb{Z}^{0}(p)$ and $0\leqslant s<p$, let $M(a,s)$ be the minimal non-negative integer such that $p^{M(a,s)}D_{an}\in\mathcal{L}(1+s)_{W_{M}(k)}$.
@@
Clearly, $M(a,0)=0$ and for all $n$, $p^{m}D_{0}\in\mathcal{L}(1)_{W_{M}(k)}\setminus\mathcal{L}(2)_{W_{M}(k)}$. In addition, if $a/v_{0}\geqslant U(p-1,0)=p-1$ then $D_{an}\in\mathcal{L}(p)_{W_{M}(k)}$. Therefore, we can assume below that $s\geqslant 1$ and $a<pv_{0}$.
Suppose $a/v_{0}=\sum_{i\geqslant 0}\theta_{i}(a)/p^{i}$ is the standard decomposition with the $p$-digits $0\leqslant\theta_{i}(a)<p$. Let $u(a)$ be the minimal index such that $\theta_{u(a)}(a)\neq 0$.
Set $u=u(a)$, $\kappa_{0}=\theta_{u}(a)$ and $\kappa_{1}=\theta_{u+1}(a)$. Note that
$\kappa_{0}/p^{u}+\kappa_{1}/p^{u+1}\leqslant a/v_{0}<\kappa_{0}/p^{u}+(\kappa_{1}+1)/p^{u+1}\,.$
Consider the relation $R_{a}(\alpha_{0},\alpha_{1},j)$ from Prop.1.2. Then
@@
and $p^{\alpha_{0}j+\alpha_{1}(j+1)}D_{an}\in\mathcal{L}(1+\alpha_{0}+\alpha_{1})_{W_{M}(k)}\,.$
Note that $\alpha_{0}+\alpha_{1}\geqslant p-1$ implies $p^{\alpha_{0}j+\alpha_{1}(j+1)}D_{an}\in\mathcal{L}(p)_{W_{M}(k)}$. If we want to obtain the relations for the ideals $\mathcal{L}(s)$ modulo $\mathcal{L}(p)$ then it makes sense to restrict ourselves with the cases where $\alpha_{0}+\alpha_{1}<p$. In particular, $\alpha_{0}/p^{j}+\alpha_{1}/p^{j+1}$ will appear as a standard decomposition in base $p$ with $p$-digits $\alpha_{0},\alpha_{1}$. Now inequality (1.3) implies one of the following options:
@@
Proceeding along these lines we obtain the following list of relations of the form $p^{M(a,s)}D_{an}\in\mathcal{L}(1+s)_{W_{M}(k)}$:
@@
For $(\alpha_{0},\alpha_{1},u)$ such that:
a) $1\leqslant\alpha_{0}\leqslant\kappa_{0}$ and $\alpha_{1}=0$, we have $s=\alpha_{0}$ and
$M(a,s)=\alpha_{0}u\,;$
b) $\alpha_{0}=\kappa_{0}$ and $1\leqslant\alpha_{1}\leqslant\kappa_{1}$, we have $s=\kappa_{0}+\alpha_{1}$ and
$M(a,s)=M(a,\kappa_{0})+\alpha_{1}(u+1)\,;$
c) $\alpha_{0}=\kappa_{0}-1$, $\alpha_{1}=\kappa_{1}+2$, we have $s=\kappa_{0}+\kappa_{1}+1$ and
$M(a,s)=M(a,\kappa_{0}+\kappa_{1})+u+2\,;$
d) $\alpha_{0}=\kappa_{0}-1$ and $\alpha_{1}=\kappa_{1}+2+t<p$ with $t\geqslant 1$, we have
$s=\kappa_{0}+\alpha_{1}$ and
$M(a,s)=M(a,\kappa_{0}+\kappa_{1}+1)+t(u+1)\,.$
@@
Taking into account that we need relations modulo $\mathcal{L}(p)_{k}$:
– for $s\in[0,\kappa_{0}]$, $M(a,s)$ is increasing in $s$ with step $u$;
– for $s\in[\kappa_{0},\min\{\kappa_{0}+\kappa_{1},p-1\}]$ the step is $u+1$;
– for $s\in[\kappa_{0}+\kappa_{1},\min\{\kappa_{0}+\kappa_{1}+1,p-1\}]$ the step is $u+2$;
– if $s\in[\kappa_{0}+\kappa_{1}+1,p-1]$ the step is $u+1$.
Proposition 1.4.
Relations from Proposition 1.3 with $s<p$ are all independent relations of the form $p^{A}D_{an}\in\mathcal{L}(1+s)_{W_{M}(k)}$, where $A\geqslant 0$ and $10\leqslant s\leqslant p-$ 1.
Proof.
The independence is obvious because for each $a$, the numbers $M(a,s)$ are strictly increasing in $s$.
Let $a/v_{0}\geqslant s_{0}/p^{j}+s_{1}/p^{j+1}$ with integers $j\geqslant 0$, $s_{0}\geqslant 1$ and $s_{1}\geqslant 0$. Note that $s_{0}$ and $s_{1}$ are not necessarily $p$-digits.
Prove that the corresponding condition $p^{m}D_{an}\in\mathcal{L}(1+s)_{W_{M}(k)}$, where $m=s_{0}j+s_{1}(j+1)$ and $s=\min\{s_{0}+s_{1},p-1\}$, is implied by one of relations listed in Proposition 1.3.
Otherwise, there is at least one triple $(\alpha_{0},\alpha_{1},u)$ listed in Prop. 1.3 such that the following two conditions hold
(1.4) $\alpha_{0}+\alpha_{1}\leqslant s_{0}+s_{1}$
(1.5) $\alpha_{0}u+\alpha_{1}(u+1)>s_{0}j+s_{1}(j+1)\,.$
Then $s_{0}\geqslant s_{0}-\alpha_{0}\geqslant(\alpha_{0}+\alpha_{1}-s_{0}-s_{1})(u+1)+(s_{0}-\alpha_{0})=\alpha_{0}u+\alpha_{1}(u+1)-(s_{0}u+s_{1}(u+1))>(s_{0}+s_{1})(j-u)$ implies that $j\leqslant u$.
If $j<u$ then $p/p^{u}>a/v_{0}\geqslant s_{0}/p^{j}$ implies $p>s_{0}p^{u-j}\geqslant p$. The contradiction.
Therefore, $j=u$. Then (1.4) implies $\alpha_{1}>s_{1}$ because
$0\geqslant(\alpha_{0}+\alpha_{1}-s_{0}-s_{1})u>s_{1}-\alpha_{1}\,.$
In particular, $\alpha_{1}\neq 0$ and the couple $(\alpha_{0},\alpha_{1})$ does not come from condition a).
The condition $\alpha_{0}+\alpha_{1}\leqslant s_{0}+s_{1}$ also implies that $\alpha_{0}<s_{0}$. This excludes the case b) because $s_{0}\leqslant\kappa_{0}$ (use that $j=u$).
It remains to consider the cases c) and d), where $\alpha_{0}=\kappa_{0}-1$ and $s_{0}=\kappa_{0}$. Here $\alpha_{0}=\kappa_{0}-1$, $\alpha_{1}\geqslant\kappa_{1}+2$, $s_{0}=\alpha_{0}+1$, and $s_{1}\leqslant\kappa_{1}$. Then we obtain the contradiction $\alpha_{0}+\alpha_{1}\geqslant\kappa_{0}+\kappa_{1}+1>s_{0}+s_{1}$. ∎
1.4. Explicit description of ideals $\mathcal{L}(w)_{W_{M}(k)}$
Recall that the ideal $\mathcal{L}(w)_{W_{M}(k)}$ is generated as $W_{M}(k)$-module by all commutators

### v5

@@ 1. Lie algebra $\mathcal{L}$ and its filtration
In this section $k\simeq\mathbb{F}_{p^{N_{0}}}$, $W(k)$ is the corresponding ring of Witt vectors, $\sigma$ is the morphism of taking $p$-th powers in $k$. The corresponding morphism $W(\sigma)$ we will denote for simplicity just by $\sigma$.
@@ 1.1. Lie algebra $\mathcal{L}$
Set $\mathbb{Z}^{+}(p)=\{a\in\mathbb{N}\ |\ \mathrm{gcd}\,(a,p)=1\}$ and $\mathbb{Z}^{0}(p)=\mathbb{Z}^{+}(p)\cup\{0\}$. Let ^$\widehat{\mathcal{L}}_{W(k)}$ be a free Lie $W(k)$-algebra with the module of (free) generators
@@
Define $\sigma$-action on ^$\widehat{\mathcal{L}}_{W(k)}$ by setting $\sigma:D_{an}\mapsto D_{a,n+1}$ and $\sigma:D_{0}\mapsto D_{0}$. Then ^$\widehat{\mathcal{L}}_{W(k)}|_{\sigma=\mathrm{id}}:=\widehat{\mathcal{L}}$ is a free Lie $\mathbb{Z}_{p}$-algebra such that ^$\widehat{\mathcal{L}}\otimes W(k)=\widehat{\mathcal{L}}_{W(k)}$. We set :=^$\mathcal{L}:=\widehat{\mathcal{L}}/C_{p}(\widehat{\mathcal{L}})\otimes\mathbb{Z}/p^{M}$. This is the maximal quotient of $\widehat{\mathcal{L}}$ of nilpotent class $<p$ and period $p^{M}$.
@@ 1.2. Ideals $\mathcal{L}(w)$, $w\geqslant 1$
Note that directly from definitions it follows that:
$U(s,0)=s$ and for $s\geqslant 1$, $U(s,m)>0$;
– if $s^{\prime}>s$ then for any $m\geqslant 1$, $U(s^{\prime},m)>U(s,m)$;
– if $m^{\prime}>m$ and $s\geqslant 1$ then $U(s,m^{\prime})<U(s,m)$.
@@
b) For $s\geqslant 1$, let $m=su+s_{1}$, where $0\leqslant s_{1}<s$. Then
$U(s,m)=s_{0}/p^{u}+s_{1}/p^{u+1}\,\text{with }s_{0}=s-s_{1}\,.$
In particular, $u,s_{0},s_{1}$ are uniquely determined by $m$ and $s$.
c) For any $m^{0},s,s^{\prime}\geqslant 0$,
$\min\{U(s,m)+U(s^{\prime},m^{\prime})\ |\ m+m^{\prime}=m^{0}\}=U(s+s^{\prime},m^{0})\,.$
@@
Therefore, the minimal value $U(s,m)$ can appear only if there is $u\in\mathbb{N}$ such that among all numbers $m_{i}$, $1\leqslant j\leqslant s$, there are $s_{0}>0$ of those which are equal to $u$ and all remaining $s_{1}=s-s_{0}$ numbers are equal to $u+1$. It remains to note that $m=m_{1}+\dots+m_{s}=us_{0}+(u+1)s_{1}=su+s_{1}\,.$
c) Suppose $U(s,m)+U(s^{\prime},m^{\prime})$ is the minimal value of the LHS. We can assume that $s,s^{\prime}\geqslant 1$. Let $m=si+s_{1}$ and $m^{\prime}=s^{\prime}i^{\prime}+s_{1}^{\prime}$, where $0\leqslant s_{1}<s$ and $0\leqslant s_{1}^{\prime}<s^{\prime}$. Then for $s_{0}=s-s_{1}$ and $s_{0}^{\prime}=s^{\prime}-s_{1}^{\prime}$, $U(s,m)=s_{0}/p^{i}+s_{1}/p^{i+1}$ and $U(s^{\prime},m^{\prime})=s_{0}^{\prime}/p^{i^{\prime}}+s_{1}^{\prime}/p^{i^{\prime}+1}$.
@@
Definition.
Let $\{\mathcal{L}(w)\}_{w\in\mathbb{N}}$ be the minimal (decreasing) filtration of ideals in $\mathcal{L}$ such that:
1) $\mathcal{L}(1)=\mathcal{L}$ and $\forall w_{1},w_{2}$, $[\mathcal{L}(w_{1}),\mathcal{L}(w_{2})]\subset\mathcal{L}(w_{1}+w_{2})$;
2) if $a/v_{0}\geqslant U(w-1,m)$ then $p^{m}D_{an}\in\mathcal{L}(w)_{W(k)}$.
Remark.
Let $\hat{\mathcal{A}}_{W(k)}$ be the enveloping algebra of ^$\hat{\mathcal{L}}_{W(k)}$. This is free associative $W(k)$-algebra. It has (as a $W(k)$-module) a $W(k)$-basis consisted of all finite products of elements from the set (1.1). We can introduce a minimal decreasing filtration of ideals $\{\hat{J}(w)\}_{w\in\mathbb{N}}$ in $\hat{\mathcal{A}}_{W(k)}$ such that
1) $\hat{J}(1)$ is the augmentation ideal of $\hat{\mathcal{A}}_{W(k)}$ and for all $w_{1},w_{2}$,
$\hat{J}(w_{1})\hat{J}(w_{2})\subset\hat{J}(w_{1}+w_{2});$
2) for all $D_{an}$ and $m\geqslant 0$, if $p^{m}a\geqslant v_{0}U(w-1,m)$ then $p^{m}D_{an}\in\hat{J}(w)\,.$
It can be easily verified that we have the induced filtration on the Lie algebra $\hat{\mathcal{L}}$ which induces our filtration $\{\mathcal{L}(w)\}_{w\in\mathbb{N}}$ on $\mathcal{L}$.
Example. Suppose $m=0$ and $s\geqslant 1$. Then $u_{0}=s_{1}=0$ and for all $w\geqslant 1$, $U(w-1,0)=w-1$. If $(w-1)v_{0}\leqslant a<wv_{0}$ then $D_{an}\in\mathcal{L}(w)_{W(k)}$. In particular, if $a\geqslant(p-1)v_{0}$ then $D_{an}\in\mathcal{L}(p)_{W(k)}$. In the modulo $p$ situation from [12] the ideals $\mathcal{L}(w)$ coincide with the ideals introduced in that paper (and denoted by $\mathcal{L}(s)$).
Part b) of above proposition 1.1 implies the following property.
@@
In the above definition of the filtration $\{\mathcal{L}(w)\}_{w\in\mathbb{N}}$ property 2) can be replaced by the following one:
$2^{\prime})$ for all $a\in\mathbb{Z}^{0}(p)$ and integers $\alpha_{0},\alpha_{1},j\geqslant 0$, it holds
@@
$p^{\alpha_{0}j+\alpha_{1}(j+1)}D_{an}\in\mathcal{L}(1+\alpha_{0}+\alpha_{1})_{W(k)}\,.$
Remark.
For $j\geqslant 1$, the relations $R_{a}(\alpha_{0},0,j)$ and $R_{a}(0,\alpha_{0},j-1)$ coincide.
1.3. Relations $p^{M(a,s)}D_{an}\in\mathcal{L}(1+s)_{W(k)}$, $s<p$
@@
Given $a\in\mathbb{Z}^{0}(p)$ and $0\leqslant s<p$, let $M(a,s)$ be the minimal non-negative integer such that $p^{M(a,s)}D_{an}\in\mathcal{L}(1+s)_{W(k)}$.
@@
Clearly, $M(a,0)=0$ and for all $m$, $p^{m}D_{0}\notin\mathcal{L}(2)_{W(k)}$. In addition, if $a/v_{0}\geqslant U(p-1,0)=p-1$ then $D_{an}\in\mathcal{L}(p)_{W(k)}$. Therefore, planning to work later modulo $\mathcal{L}(p)_{W(k)}$ we can assume below that $s\geqslant 1$ and $1\leqslant a<(p-1)v_{0}$.
Introduce the integers $u(a)\geqslant 0$, $1\leqslant\kappa_{0}(a)<p$ and $0\leqslant\kappa_{1}(a)<p$ such that
$\kappa_{0}/p^{u}+\kappa_{1}/p^{u+1}\leqslant a/v_{0}<\kappa_{0}/p^{u}+(\kappa_{1}+1)/p^{u+1}\,,$
where we set $u=u(a)$, $\kappa_{0}=\kappa_{0}(a)$ and $\kappa_{1}=\kappa_{1}(a)$ to simplify the notation. These integers are uniquely determined by $a<(p-1)v_{0}$.
Now relation $R_{a}(\alpha_{0},\alpha_{1},j)$ from Prop.1.2 means that
@@
implies $p^{\alpha_{0}j+\alpha_{1}(j+1)}D_{an}\in\mathcal{L}(1+\alpha_{0}+\alpha_{1})_{W(k)}\,.$
Note that $\alpha_{0}+\alpha_{1}\geqslant p-1$ implies $p^{\alpha_{0}j+\alpha_{1}(j+1)}D_{an}\in\mathcal{L}(p)_{W(k)}$. If we want to obtain basic relations for the ideals $\mathcal{L}(s)$ modulo $\mathcal{L}(p)$ then it makes sense to restrict ourselves with the cases where $\alpha_{0}+\alpha_{1}<p$. In particular, $\alpha_{0}/p^{j}+\alpha_{1}/p^{j+1}$ will appear as a standard decomposition in base $p$ with $p$-digits $\alpha_{0},\alpha_{1}$. Now inequality (1.3) is equivalent to one of the following options:
@@
Proceeding along these lines introduce for a given $a\in\mathbb{Z}^{+}(p)$, the following sets of “basic” relations $R_{a}(\beta_{0},\beta_{1},u)$ (where $u=u(a)$) which provide the condition $p^{M(a,s)}D_{an}\in\mathcal{L}(1+s)_{W(k)}$:
[…]
$s=\kappa_{0}+\beta_{1}$, $M(a,s)=M(a,\kappa_{0})+\beta_{1}(u+1)$;
3) $\mathcal{R}^{III}_{a}=\{R_{a}(\kappa_{0}-1,\kappa_{1}+2,u)\ |\ \kappa_{0}+\kappa_{1}\leqslant p-2\}$; here $s=\kappa_{0}+\kappa_{1}+1$ and $M(a,s)=M(a,\kappa_{0}+\kappa_{1})+u+2$;
4) $\mathcal{R}^{IV}_{a}=\{R_{a}(\kappa_{0}-1,\kappa_{1}+2+t,u)\ |\ t\geqslant 1,\kappa_{0}+\kappa_{1}+1+t\leqslant p-1\}$; here $s=\kappa_{0}+\kappa_{1}+1+t$ and $M(a,s)=M(a,\kappa_{0}+\kappa_{1}+1)+t(u+1)$.
@@
Given $a\in\mathbb{Z}^{+}(p)$ the above listed relations imply all relations $R_{a}(\alpha_{0},\alpha_{1},j)$ modulo $\mathcal{L}(p)_{W(k)}$.
Proof.
Consider the relation $R_{a}(\alpha_{0},\alpha,j)$, i.e. it holds $a/v_{0}\geqslant\alpha_{0}/p^{j}+\alpha_{1}/p^{j+1}$ with integers $j\geqslant 0$, $\alpha_{0},\alpha_{1}\geqslant 0$. Note that $\alpha_{0}$ and $\alpha_{1}$ are not necessarily $p$-digits. Prove that the corresponding condition $p^{m}D_{an}\in\mathcal{L}(1+s)_{W(k)}$, where $m=\alpha_{0}j+\alpha_{1}(j+1)$ and $s=\min\{\alpha_{0}+\alpha_{1},p-1\}$, is implied by one of the relations $R_{a}(\beta_{0},\beta_{1},u)$ from the sets in above 1)-4).
Otherwise, there is at least one relation $R_{a}(\beta_{0},\beta_{1},u)$ listed in 1)-4), such that the following two conditions hold
(1.4) $\beta_{0}+\beta_{1}\leqslant\min\{\alpha_{0}+\alpha_{1},p-1\}$
(1.5) $\beta_{0}u+\beta_{1}(u+1)>\alpha_{0}j+\alpha_{1}(j+1)\,.$
Then $\alpha_{0}\geqslant\alpha_{0}-\beta_{0}\geqslant(\beta_{0}+\beta_{1}-\alpha_{0}-\alpha_{1})(u+1)+(\alpha_{0}-\beta_{0})=\beta_{0}u+\beta_{1}(u+1)-(\alpha_{0}u+\alpha_{1}(u+1))>(\alpha_{0}+\alpha_{1})(j-u)$ implies that $j\leqslant u$.
$\bullet$ Suppose $\alpha_{0}\neq 0$.
If $j<u$ then $p/p^{u}>a/v_{0}\geqslant\alpha_{0}/p^{j}$ implies $p>\alpha_{0}p^{u-j}\geqslant p$. The contradiction. Therefore, $j=u$. Then (1.5) implies $\beta_{1}>\alpha_{1}$ because
$0\geqslant(\beta_{0}+\beta_{1}-\alpha_{0}-\alpha_{1})u>\alpha_{1}-\beta_{1}\,.$
In particular, $\beta_{1}\neq 0$ and $R_{a}(\beta_{0},\beta_{1},u)\notin\mathcal{R}_{a}^{I}$. Now condition (1.4) implies that $\beta_{0}<\alpha_{0}$. This excludes the set $\mathcal{R}_{a}^{II}$ because $\alpha_{0}\leqslant\kappa_{0}$ (use that $j=u$).
It remains to consider the sets $\mathcal{R}_{a}^{III}$ and $\mathcal{R}_{a}^{IV}$, where $\beta_{0}=\kappa_{0}-1$ and $\alpha_{0}=\kappa_{0}$. Here $\beta_{1}\geqslant\kappa_{1}+2$ and $\alpha_{1}\leqslant\kappa_{1}$. This gives the contradiction $\beta_{0}+\beta_{1}\geqslant\kappa_{0}+\kappa_{1}+1>\alpha_{0}+\alpha_{1}$.
$\bullet$ Suppose $\alpha_{0}=0$ and $\alpha_{1}\neq 0$.
Proceed similarly to the above case where $\alpha_{0}\neq 0$. Here $0>\alpha_{1}(j-u)$ implies $j+1\leqslant u$. If $j+1<u$ we obtain the following contradiction $p/p^{u}>a/v_{0}\geqslant\alpha_{1}/p^{j+1}$, i.e. $p>\alpha_{1}p^{u-j-1}\geqslant p$. So, $u=j+1$.
Now $a/v_{0}\geqslant\alpha_{1}/p^{u}$ implies that $\beta_{0}\geqslant\alpha_{1}$ and condition 1.4 implies $\beta_{1}=0$. Therefore, $R_{a}(\alpha_{0},\alpha_{1},j)=R_{a}(\alpha_{1},0,u)\in\mathcal{R}_{a}^{I}$. ∎
@@
Given $a\in\mathbb{Z}^{+}(p)$ the relations from 1)-4) are independent:
a) if $u(a)=0$ then relations $R_{a}(\beta_{0},0,0)\in\mathcal{R}_{a}^{I}$ could be replaced by the relation $R_{a}(\kappa_{0},0,0)$. This relation together with remaining ones $R_{a}(\beta_{0},\beta_{1},0)\in\mathcal{R}_{a}^{I}$ are independent because the appropriate values $M(a,s)$ strictly increase in $s$.
b) if $u(a)\neq 0$ then (for a fixed $a$) all $M(a,s)$ strictly increase together with $s$ and, therefore, are independent.
1.4. Explicit description of ideals $\mathcal{L}(w)_{W(k)}$
Recall that the ideal $\mathcal{L}(w)_{W(k)}$ is generated as $W(k)$-module by all commutators